OpenStudy (anonymous):

erasers cost $5 per carton and pencils cost $7 per carton. If an order comes in for a total of 15 cartons for $85, how many cartons of each were bought?

OpenStudy (anonymous):

let x cartons were bought for eraser and y cartons were bought for pencils then by the problem x+y=15 5x+7y=85

OpenStudy (anonymous):

multiplying first eq by 5 we get 5x+5y=75 5x+7y=85 (-) 5x+5y=75 ______________________(subtracting ) 2y=10 hence y=10/2 or y=5 now x+y=15 x=15-y=15-5=10 10 cartoons of eraser were bought and 5 cartoons of pencils were bought